# Dual axis atomic magnetometer and gyroscope enabled by nuclear spin perturbation

### v2

@@ VI Conclusion
Through the applications of magnetic and optical pulses, we have devised a new protocol for extracting signals from an atomic comagnetometer. The scheme is able to measure magnetic and non-magnetic effects on two orthogonal axes with a noise performance that, in ideal conditions, should be comparable to the state-of-the-art continuous wave scheme, where only a single parameter is measured. A proof-of-principle experimental realisation based on ${}^{87}\mathrm{Rb}$ and ${}^{129}\mathrm{Xe}$ was presented with promising results. In particular, we demonstrated that the PANCo scheme effectively suppresses cross-talk between magnetic and non-magnetic signals to a level that is comparable to the best self-compensated alkali-noble comagnetometers. This result was achieved with minimal optimisation, yet the scheme has a wide parameter space for performance gains in terms of the applied pulses, the applied bias fields, gas mixtures and investigations of techniques to diagnose and minimise other technical noise. In future work we plan to explore these factors to uncover the practical limits of the scheme.

### v3

@@ VI Conclusion
Through the applications of magnetic and optical pulses, we have devised a new protocol for extracting signals from an atomic comagnetometer. The scheme is able to measure magnetic and non-magnetic effects on two orthogonal axes with a noise that, in ideal conditions, should be comparable to the state-of-the-art continuous wave scheme, where only a single parameter is measured. A proof-of-principle experimental realisation based on ${}^{87}\mathrm{Rb}$ and ${}^{129}\mathrm{Xe}$ was presented with promising results. In particular, we demonstrated that the PANCo scheme effectively suppresses cross-talk between magnetic and non-magnetic signals to a level that is comparable to the best self-compensated alkali-noble comagnetometers. This result was achieved with minimal optimisation, yet the scheme has a wide parameter space for performance gains in terms of the applied pulses, the applied bias fields, gas mixtures and investigations of techniques to diagnose and minimise other technical noise. In future work we plan to explore these factors to uncover the practical limits of the scheme.